THE FIRST APPLICATION OF DOPPLER-FREE LASER POLARIZATION AND OPTICAL-OPTICAL DOUBLE RESONANCE POLARIZATION LABELING SPECTROSCOPIES TO A LARGE MOLECULE

M. H. KABIR, S. KASAHARA, Y. TATAMITANI and \underline{HAJIME KATO}, Molecular Photoscience Research Center, Kobe University, Kobe 657-8501, Japan; W. DEMTRÖDER, Fachbereich Physik, Universität Kaiserslautern, D-67663 Kaiserslautern, Germany; M. BABA, Faculty of Integrated Human Studies, Kyoto University, Kyoto 606-8501, Japan.

For the first time Doppler-free laser polarization (DFLP) spectroscopy is successfully applied to a large polyatomic molecule. Rotationally resolved spectra of the 33_{0}^{1} vibronic bands of the Ã ^{1}B_{1u} ← X̃ ^{1}A_{g} electronic transition of gas phase naphthalene at room temperature have been measured with the techniques of DFLP spectroscopy. The typical width of the observed spectral lines was 15 MHz, and the absolute wave numbers were measured with an accuracy of better than ±0.0002 cm^{−1}. In order to assign the dense and complicated parts of the spectra, the technique of Doppler-free opticaloptical double resonance polarization labeling (DFOOPL) spectroscopy was used. Both V-type and Λ-type transitions were observed, and these signals were found to be very useful for the unambiguous assignment of the complicated lines of DFLP spectrum. The rotational constants A, B, and C, the symmetric-top centrifugal distortion constants Delta_{J}, Δ_{JK}, and Δ_{K} and two asymmetric-top distortion constants δ_{J} and δ_{K} were determined by a least-squares fitting of over 4000 spectral lines in the range J = 9 − 150 and K_{a} = 0 − 37 with a standard deviation of 0.0002 cm^{−1}.